OpenStudy (anonymous):

The length of a rectangle is increasing at a rate of 3 cm/s and its width is increasing at a rate of 9 cm/s. When the length is 10 cm and the width is 5 cm, how fast is the area of the rectangle increasing?

OpenStudy (anonymous):

Well, you want to use \[ A = l\times w \]

OpenStudy (anonymous):

And then you differentiate.

OpenStudy (anonymous):

I would like to see it step by step if possible

OpenStudy (anonymous):

same idea as before Area of the rectangle is \(A=xy\) taking the derivative, and using the product rule, you get \[A'=xy'+yx'\] plug in the numbers your are given

OpenStudy (anonymous):

\[ \frac{dA}{dt} = \frac{dl}{dt}w + l\frac{dw}{dt} \]

OpenStudy (anonymous):

We are given that \(dl/dt = 3\text{cm/s}\) and \(dw/dt = 9\text{cm/s}\)

OpenStudy (anonymous):

Substituting that in, we get: \[ \frac{dA}{dt} = 3w\ \text{cm/s} + 9l\ \text{cm/s} \]

OpenStudy (anonymous):

They also give us \(l\) and \(w\), so we have to substitute them in too.

OpenStudy (anonymous):

what about just being in cm

OpenStudy (anonymous):

never mind, i was trying to figure the wrong thing out that was in the question itself